# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ II.3 The baseline model as a function of inducer concentrations
Suppose now that each repressor produced from the three-gene toggle circuit in Fig. 1 can be controlled by its own inducer, each at concentrations $c_{1}$, $c_{2}$, and $c_{3}$. We will make the reasonable assumption that each repressor responds to induction with the same sensitivity and thus by the same probability function. As shown in Fig. 2(B), this means that expression of repressor $1$ depends on the concentrations of active repressors 22 and $3$, namely, $p_{\text{act}}(c_{2})\times R_{2}$ and $p_{\text{act}}(c_{3})\times R_{3}$. We similarly scale the regulatory contribution of $R_{1}$ to expression of $R_{2}$ and $R_{3}$ by the probability $p_{\text{act}}(c_{1})$.
Eqns. 24 describing this system’s dynamics thus become
@@
We can simplify our model to a dimensionless form through several assumptions. First, we assume for analytic convenience that all repressors are produced (i.e., translated) and degraded at essentially the same maximal rates, independent of the underlying dynamics of transcription. This sets $a_{1}=a_{2}=a_{3}=a$ and $\tau_{1}=\tau_{2}=\tau_{3}=\tau$. We will also assume that a given repressor binds to different promoter sites with the same affinity, such that $K_{21}=K_{31}\equiv K_{1}$, $K_{12}=K_{32}\equiv K_{2}$, and $K_{13}=K_{23}\equiv K_{3}$. This allows us to re-express Eqns. 1416 by transforming $\bar{t}=t/\tau$, $\bar{R}_{i}=R_{i}/K_{i}$, and $\bar{a}_{i}=\tau a/K_{i}$, from which we obtain the dimensionless equations
@@
Finally, for most of the discussion in this work we will simplify to a symmetric system in which all repressors bind at all sites with equal affinity such that $K_{1}=K_{2}=K_{3}=K$ (we will examine how breaking this symmetry affects shifts in dynamics in later sections). This assumption then means that $\bar{a}_{1}=\bar{a}_{2}=\bar{a}_{3}=\bar{a}$, and we arrive at the system of equations
@@
The original model defined by Eqns. 24 thus transforms from a high-dimensional parameter space as a function of inherent system properties to an at most three-dimensional parameter space defined by inducer concentrations $c_{1}$, $c_{2}$, and $c_{3}$ that can be directly controlled in experimental settings. Note that the simplifying assumptions introduced here are in no sense necessary and we could examine the much less symmetrical situations as well, but we find that this parameter choice makes the underlying dynamics most transparent.
@@ III Regimes of multistability
Figure 4: Heatmaps tracking the bifurcation threshold beyond which the three-gene toggle switch can no longer have an $\bar{R}_{1}$-dominant steady state. Note that we have defined the limit steady state case here as satisfying $\bar{R}_{2}^{n}$, $\bar{R}_{3}^{n}\rightarrow 0$, with $\bar{R}_{2}=\bar{R}_{3}\leq\varepsilon$. The results shown are specific to the choice $\varepsilon=\sqrt[n]{0.01}$, i.e., the boundary for allowing tristable dynamics to be possible (A) The threshold probability of active repressor as a function of Hill coefficient $n$ and mRNA expression rate $\bar{a}$. Note that the beige shaded region for smaller axis values denotes a regime in which the steady state is never possible at any inducer concentration. (B) The threshold in terms of inducer concentration for the MWC model as specified in Fig. 3(A).
For a given Hill exponent $n$ and maximal (dimensionless) production rate $\bar{a}$, we now explore the dynamic profiles possible for the three-gene toggle switch as a function of inducer concentrations. In particular, we quantify the thresholds, or bifurcations, as a function of inducer that bring about fundamental shifts in dynamics. Our graphical analysis throughout this section will consider the large cooperativity case $n=4$ and choose $\bar{a}=2$. While we will also address how much the dynamic thresholds change with parameter choices, our focus in this section is specifically on the effects of induction at various levels. To this end, we begin by considering a system controlled only by a single inducer.
@@ III.1 Single inducer
Suppose a single inducer controls the dynamics of the three-gene toggle switch by regulating the activity of $R_{1}$. This means that $p_{\text{act}}(c_{1})\equiv p_{\text{act}}(c)$ and $p_{\text{act}}(c_{2})=p_{\text{act}}(c_{3})=1$ such that $\bar{R}_{2}$ and $\bar{R}_{3}$ are always maximally active. Substituting these definitions into Eqns. 2022, we obtain the single-inducer model
@@
For a given inducer concentration $c$, one then sets the above equations to zero and solves to obtain the possible steady state expressions.
We can first gain some analytical insight from these equations by considering certain limit cases, specifically those regarding the existence of single-repressor-dominant steady states, as a function of inducer concentration.
Consider the existence of an $\bar{R}_{2}$-dominant steady state whereby $\bar{R}_{1}=\bar{R}_{3}=\varepsilon$ such that $\bar{R}_{1}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$. It follows from setting Eqn. 24 to zero that such a steady-state requires $\bar{R}_{2}=\bar{a}$. By similarly setting Eqns. 23 and 25 to zero, it follows that $\bar{R}_{1,\>3}=\bar{a}/(1+\bar{R}_{2}^{n})$. We can therefore conclude that the system always has the following steady state regardless of inducer concentration:
$\bar{R}_{ss}=(\bar{R}_{1},\>\bar{R}_{2},\>\bar{R}_{3})=\Big(\frac{\bar{a}}{1+\bar{a}^{n}},\bar{a},\frac{\bar{a}}{1+\bar{a}^{n}}\Big)$ (26)
Similar logic for the existence of an $\bar{R}_{3}$-dominant state leads to a similar predicted steady state
@@
The conditions for an $\bar{R}_{1}$-dominant steady state, where $\bar{R}_{2}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$, follow from the same logic in Eqns. 2325 as $\bar{R}_{1}=\bar{a}$ and
@@
Unlike the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant states in Eqns. 26 and 27, Eqn. 28 indicates that for the $\bar{R}_{1}$-dominant steady state the concentrations of non-induced repressors $\bar{R}_{2}$ and $\bar{R}_{3}$ depend on inducer concentration $c$. As $c$ increases and the probability of activity decreases, the corresponding repressor concentrations at steady state increase. This also means, however, that at a sufficiently high inducer concentration, $\bar{R}_{2}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$ can no longer hold true and the steady state can no longer exist. Assuming that $\bar{R}_{2}=\bar{R}_{3}\leq\epsilon$ satisfies the specified condition, we can derive the inducer bifurcation threshold by substituting this inequality into Eqn. 28 and rearranging, leading to the result
@@
At equality, the probability of Eqn. 29 thus represents a bifurcation threshold, where for lower probabilities (i.e, higher inducer concentrations) the system can no longer exist at an $\bar{R}_{1}$-dominant steady state.
@@
Figure 5: Dynamics of the three-gene toggle switch with $n=4$ and $\bar{a}=2$ for increasing inducer concentration $c$. Allosteric regulation is defined by the probability curve shown in Fig. 3(A). (A)-(B) Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady state expression as a function of inducer concentration. (C) Fixed points for the three-gene toggle switch at a low inducer concentration. (D) Fixed points for the three-gene toggle switch at an intermediate inducer concentration. (E) Fixed points for the three-gene toggle switch at a high inducer concentration. In (C)-(E) the stability of each fixed point is color-coded as in panels (A) and (B), with the expression levels for each fixed point colored as in Fig. 1.
@@
The results in Eqns. 2629 therefore indicate that at low inducer concentration the system is able to arrive at a steady state in which any gene amongst the three can dominate expression. When the probability of activity drops below the threshold analytically approximated in Eqn. 29, however, the presence of inducer is sufficient to suppress $\bar{R}_{1}$ activity, and the system can thus no longer maintain a steady state with $\bar{R}_{1}$-dominant expression. The sensitivity of $\bar{R}_{1}$ activity to tuning then depends on cooperativity (as approximated by Hill coefficient $n$) and on the protein expression rate $\bar{a}$.
@@
Fig. 4(A) plots the probability threshold of Eqn. 29 as a function of $n$ and $\bar{a}$. Note that the beige region of the heatmap depicts a regime where an $\bar{R}_{1}$-dominant steady state is not possible at any inducer concentration. The threshold for existence relies on sufficient cooperativity. Note, though, that when $\bar{a}\gtrsim 2$, increasing the expression rate has little effect on this cooperativity bound. Fig. 4(B) plots the same relationship between the parameters, now explicitly as a function of inducer concentration (from the probability definition of Eqn. 6 in Section II.2).
@@
The figure tracks cooperativities up to $n=10$. While it is not common for known multimeric transcription factors to extend beyond tetramers, plotting higher $n$ acknowledges the uncertainty that remains in the field regarding the cooperative mechanics of eukaryotic regulation and particularly of enhancers [64]. Indeed, there are emerging cases of eukaryotic transcription factors functioning as higher-order homomultimeric complexes, such as FOXP3 in regulatory T-cells [65]. It is therefore useful to observe whether such putative higher-order cooperativities would influence bifurcations observed in the system.
@@
We observe from the figure that the bifurcation threshold for this dynamic shift is most sensitive to parameter changes at the onset of the expression domain (i.e., at minimum required $n$ and $\bar{a}$). For $n\gtrsim 4$ and $\bar{a}\gtrsim 2$, both parameters are sufficiently high that increasing either has little effect on the threshold probability and corresponding inducer concentration. It is therefore reasonable that, to understand the complexity of dynamics possible for the toggle switch, we choose $n=4$ as representative of dynamics in the highly-cooperative regime of the system for the remainder of this work. This is also a biologically reasonable choice, given the existence of homotetrameric transcription factors such as p53 that are crucial to cell fate decisions [66, 67, 68].
@@
To show more completely how the dynamics evolve with increasing inducer concentration, we now choose a specific set of parameters (in this case from the highly cooperative regime of $n=4$ and $\bar{a}=2$) for Eqns. 2325. At each inducer concentration, we numerically solve for the fixed point repressor concentrations of these equations. We then determine the stability of each fixed point by performing a linear stability analysis. We achieve this by evaluating the Jacobian which, for our three-gene network modeled by differential equations of the form $d\bar{R}_{i}/dt=f_{i}(\bar{R}_{1},\bar{R}_{2},\bar{R}_{3})$, is a $3\times 3$ matrix with entries
@@
The Jacobian matrix provides a first-order linear approximation of behavior evaluated near a fixed point, with its eigenvectors specifying the primary directions of dynamical motion from the fixed point, and with its eigenvalues reflecting the rates at which perturbations to the fixed point grow or decay along the corresponding eigenvectors. If all eigenvalues have negative real parts, perturbation in any eigenvector direction of expression space will decay back to the fixed point, indicating it to be a stable equilibrium. If any of the eigenvalues have a positive real part, however, perturbation in the corresponding eigenvector direction(s) will grow exponentially and the system will move away from the fixed point, marking it as unstable.
@@
For the three-gene models examined throughout this study, the unstable fixed points observed are characterized as saddle points. Mathematically, this distinction arises because at least one eigenvalue is negative. Dynamic trajectories local to the fixed point thus move away in some dimensional directions (positive eigenvalues) and approach in others (negative eigenvalues). The number of positive eigenvalues determines the number of dimensional directions of “escape” from the fixed point and thus the “index” of the saddle. If we consider the dynamics of protein expression within a potential landscape, expression levels local to the highest index saddle essentially follow a hierarchical dynamic flow through expression space toward stable states, guided by the presence of saddle points with decreasing index.
@@
Fig. 5 plots the resulting changes in the number of fixed points and their expression levels through bifurcation diagrams, isolating the trajectories of the $\bar{R}_{1}$ coordinate of fixed points in Panel (A), and the (overlapping) $\bar{R}_{2}$ and $\bar{R}_{3}$ coordinate fixed point trajectories in Panel (B). Three types of fixed points emerge from analysis of the Jacobian. Stable equilibria are denoted in blue, unstable index-1 saddle points in red, and unstable index-2 saddles in yellow.
@@
At an intermediate inducer concentration, a saddle bifurcation occurs when a pair of fixed points annihilate each other. The intermediate regime, represented in Fig. 5(D), is still tristable, but the three stable states are now only connected by two saddle points. From an initial state with sufficient concentrations of all three repressors, the system can thus evolve dynamically in one of two ways: by either falling into a bistable potential well favoring $\bar{R}_{2}$ or $\bar{R}_{3}$ expression, or by following a trajectory directly stabilizing toward favoring $\bar{R}_{1}$ expression. Dynamics are therefore more restricted in this intermediate regime than in the low inducer regime, and no longer equally likely to evolve toward any given stable state.
Finally, Fig. 5(E) shows fixed point expression levels at a high inducer concentration, where the presence of inducer sufficiently suppresses the activity of $\bar{R}_{1}$, and the system undergoes a saddle bifurcation to collapse down to a bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant expression.
@@
Fig. 5 thus allows us to determine the inducer concentration thresholds at which dynamic shifts occur, including the shift between tristable and bistable dynamics. Given a distribution of stable state expression levels in the data, we could then distinguish between systems in the low inducer regime or in the more dynamically restricted intermediate regime. Fig. 5(B) also highlights that it is only at intermediate inducer concentrations that the system does not necessarily stabilize to a state exclusively favoring a single gene’s expression, which uniquely distinguishes the low vs. intermediate inducer concentration regimes of tristability.
@@
Figure 6: Change in inducer concentration range (shaded region) of the intermediate dynamic regime for increasing dissociation constant $K_{I}$, overlaid with the corresponding activity probability curves from Fig. 3. Note that regardless of $K_{I}$ parametrization the intermediate regime is found within the same range of probabilities $0.749\leq p_{\text{act}}(c)\leq 0.867$ (bounded in grey).
@@
Given the relatively narrow concentration range for the intermediate dynamic regime, it is natural to ask how sensitive this window is to changes in allosteric interaction and to changes in the cooperativity, or ultrasensitivity, of expression response to induction. Regarding sensitivity to allostery, altering the ratio of $K_{A}$ and $K_{I}$ for an inducer binding to a repressor shifts the inflection point of the probability curve for activity toward higher inducer concentrations, as previously seen in Fig. 3. We thus observe in Fig. 6 that the log scale range of inducer concentrations falling within the intermediate dynamic regime does not change but shifts toward increasing value ranges with increasing $K_{I}$. Since altering the ratio of the dissociation constants does not affect the slope of the probability curve at the inflection point, the intermediate dynamic regime (for otherwise fixed values of $m$ and $\Delta\epsilon$) exists consistently across different $K_{A}/K_{I}$ for $0.749\leq p_{\text{act}}(c)\leq 0.867$.
@@
Figure 7: Change in the regions of multistability for $n\in[3,4]$. Each dynamic region is color-coded as denoted in Fig. 5, with the seven-fixed point tristable regime (purple), five-fixed point tristable regime (light purple), and three-fixed point bistable regime (red). Note that for approximately $n\leq 3$ the most complex dynamic regime is the light purple tristable regime with five fixed points.
@@
The size of the intermediate regime may more likely change with a decrease in cooperativity. Fig.  4(B) indicates that at $\bar{a}=2$, bistability is possible for $n>3$, but the threshold for shifting from tristability to bistability shifts toward lower inducer concentrations for $n<4$. It is thus unclear from Fig. 4 alone whether the range of inducer concentrations allowing the intermediate dynamic regime would expand, contract, or remain the same for a smaller cooperative Hill coefficient. Indeed, Fig. 7 demonstrates how the regimes of multistability evolve in response to varying cooperativity. As the Hill coefficient increases from $n=3$ (the bound above which the system can exist in one of three possible dynamic regimes), the inducer concentration at which the system bifurcates from tristable to bistable dynamics rises but not significantly when viewed in the log scale. The threshold separating the two distinct regimes of tristability, however, does significantly change. The inducer concentration range corresponding to the intermediate regime diminishes with increasing cooperativity, since increased cooperativity means a more sensitive response in gene expression to the binding of repressors that emphasizes the bias away from $\bar{R}_{1}$ expression with increased induction. The tristable threshold begins to approach an apparent asymptote at $n=4$, confirming our earlier finding that increasing cooperativity to $n=4$ has little effect on the dynamics.

### v2

@@ II.3 The baseline model as a function of inducer concentrations
Suppose now that each repressor produced from the three-gene toggle circuit in Fig. 1 can be controlled by its own inducer, each at concentrations $c_{1}$, $c_{2}$, and $c_{3}$. We will make the reasonable assumption that each repressor responds to induction with the same sensitivity and thus in accord with the same probability function. As shown in Fig. 2(B), this means that expression of repressor $1$ depends on the concentrations of active repressors 22 and $3$, namely, $p_{\text{act}}(c_{2})\times R_{2}$ and $p_{\text{act}}(c_{3})\times R_{3}$. We similarly scale the regulatory contribution of $R_{1}$ to expression of $R_{2}$ and $R_{3}$ by the probability $p_{\text{act}}(c_{1})$. Eqns. 24 thus become
@@
We can simplify the model to a dimensionless form through several assumptions. First, we assume for analytic convenience that all repressors are produced and degraded at the same maximal rates, independent of the underlying dynamics of transcription. This sets $a_{1}=a_{2}=a_{3}\equiv a$ and $\tau_{1}=\tau_{2}=\tau_{3}\equiv\tau$. We also assume that a given repressor binds to different promoter sites with the same affinity, such that $K_{21}=K_{31}\equiv K_{1}$, $K_{12}=K_{32}\equiv K_{2}$, and $K_{13}=K_{23}\equiv K_{3}$. This allows us to re-express Eqns. 79 in dimensionless form, transforming $\bar{t}=t/\tau$, $\bar{R}_{i}=R_{i}/K_{i}$, and $\bar{a}_{i}=\tau a/K_{i}$ to obtain
@@
Finally, for most of the discussion we consider a symmetric system in which all repressors bind at all sites with equal affinity such that $K_{1}=K_{2}=K_{3}\equiv K$. (We will examine how breaking this symmetry affects shifts in dynamics in later sections.) This means that $\bar{a}_{1}=\bar{a}_{2}=\bar{a}_{3}=\bar{a}$, and we arrive at the system of equations
@@
The original model in Eqns. 24 thus collapses from a high-dimensional parameter space to at most a three-dimensional space defined by inducer concentrations $c_{1}$, $c_{2}$, and $c_{3}$, which can be directly controlled experimentally. Note that while the symmetry introduced here is not necessary, it makes the underlying dynamics more transparent.
@@ III Regimes of multistability
Figure 4: Heatmaps tracking the bifurcation threshold beyond which the three-gene toggle switch can no longer have an $\bar{R}_{1}$-dominant steady state and be tristable. Note that we define the limit steady state case here as satisfying $\bar{R}_{2}^{n}$, $\bar{R}_{3}^{n}\rightarrow 0$, with $\bar{R}_{2}=\bar{R}_{3}\leq\varepsilon$. The results shown are specific to the choice $\varepsilon=\sqrt[n]{0.01}$. (A) The threshold probability of active repressor as a function of Hill coefficient $n$ and mRNA expression rate $\bar{a}$. Note that the beige-shaded region for smaller axis values denotes a regime in which the steady state is never possible at any inducer concentration. (B) The threshold specified in Fig. 3(A), now in terms of inducer concentration for the MWC model.
For a given Hill exponent $n$ and maximal (dimensionless) production rate $\bar{a}$, we now explore the dynamic profiles possible for the three-gene toggle switch as a function of inducer concentrations. In particular, we quantify the thresholds, or bifurcations, as a function of inducer that bring about fundamental shifts in dynamics. Our graphical analysis in this section will consider the large cooperativity case where $n=4$, along with setting $\bar{a}=2$. While we will also examine how much the dynamic thresholds shift as we alter parameters, our focus in this section is specifically on the effects of induction at various intensities. To this end, we begin by considering a system tuned by a single inducer.
@@ III.1 Single inducer
Suppose a single inducer controls the dynamics of the three-gene toggle switch by regulating the activity of $R_{1}$. This means that $p_{\text{act}}(c_{1})\equiv p_{\text{act}}(c)$ and $p_{\text{act}}(c_{2})=p_{\text{act}}(c_{3})=1$ such that $\bar{R}_{2}$ and $\bar{R}_{3}$ are always maximally active. Substituting these definitions into Eqns. 1315, we obtain the single-inducer model
@@
For a given inducer concentration $c$, one then sets the above equations to zero and solves to obtain the steady state expressions.
We can first gain some analytical insight from the model by considering certain limit cases, specifically those regarding the existence of single-repressor-dominant steady states, as a function of inducer concentration. Consider the existence of an $\bar{R}_{2}$-dominant steady state whereby $\bar{R}_{1}=\bar{R}_{3}=\varepsilon$ such that $\bar{R}_{1}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$. It follows from setting Eqn. 17 to zero that such a steady-state requires $\bar{R}_{2}=\bar{a}$. By similarly setting Eqns. 16 and 18 to zero, it also follows that $\bar{R}_{1,3}=\bar{a}/(1+\bar{R}_{2}^{n})$. We therefore conclude that the system always has the following steady state regardless of inducer concentration:
$\bar{R}_{ss}=(\bar{R}_{1},\>\bar{R}_{2},\>\bar{R}_{3})=\Big(\frac{\bar{a}}{1+\bar{a}^{n}},\bar{a},\frac{\bar{a}}{1+\bar{a}^{n}}\Big).$ (19)
Similar logic for the existence of an $\bar{R}_{3}$-dominant state leads to a corresponding predicted steady state
@@
The conditions for an $\bar{R}_{1}$-dominant steady state, where $\bar{R}_{2}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$, follow from the same logic in Eqns. 1618 as $\bar{R}_{1}=\bar{a}$ and
@@
Unlike the $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant states in Eqns. 19 and 20, however, Eqn. 21 indicates that for the $\bar{R}_{1}$-dominant steady state the concentrations of non-induced repressors $\bar{R}_{2}$ and $\bar{R}_{3}$ depend on inducer concentration $c$. As $c$ increases and the probability of activity decreases, the corresponding repressor concentrations at steady state rise. This also means that at a sufficiently high inducer concentration, $\bar{R}_{2}^{n},\>\bar{R}_{3}^{n}\rightarrow 0$ no longer holds and the steady state can no longer exist. Assuming $\bar{R}_{2}=\bar{R}_{3}\leq\epsilon$ satisfies the specified condition, we derive the inducer bifurcation threshold by substituting this inequality into Eqn. 21 and rearranging to obtain the result
@@
At equality, the probability in Eqn. 22 represents a bifurcation threshold, where for lower probabilities (i.e, higher inducer concentrations) the system can no longer exist at an $\bar{R}_{1}$-dominant steady state.
@@
Figure 5: Dynamics of the three-gene toggle switch with $n=4$ and $\bar{a}=2$ for increasing inducer concentration $c$. Allosteric regulation is defined by the probability curve shown in Fig. 3(A). (A)-(B) Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady state expression as a function of inducer concentration. (C) Fixed points for the three-gene toggle switch at a low inducer concentration. (D) Fixed points for the three-gene toggle switch at an intermediate inducer concentration. (E) Fixed points for the three-gene toggle switch at a high inducer concentration. In (C)-(E) the stability of each fixed point is color-coded as in panels (A) and (B), and each fixed point is labeled with its corresponding expression levels, colored as in Fig. 1.
@@
Eqns. 1922 therefore indicate that at low inducer concentration the system can arrive at a steady state in which any gene among the three can dominate expression. When the probability of activity drops below the threshold in Eqn. 22, however, the presence of inducer is sufficient to suppress $\bar{R}_{1}$ activity, and the system can no longer maintain a steady state with $\bar{R}_{1}$-dominant expression. The sensitivity of $\bar{R}_{1}$ activity to tuning then depends on cooperativity (as approximated by Hill coefficient $n$) and on the protein expression rate $\bar{a}$.
@@
Fig. 4(A) plots the probability threshold in Eqn. 22 as a function of $n$ and $\bar{a}$. Note that the beige region of this heatmap depicts a regime where an $\bar{R}_{1}$-dominant steady state is not possible at any inducer concentration. When $\bar{a}\gtrsim 2$, increasing the expression rate has little effect on this cooperativity bound. Fig. 4(B) plots the same relationship between the parameters, now explicitly as a function of inducer concentration (from the probability definition of Eqn. 6 in Section II.2).
@@
Fig. 4 tracks cooperativities up to $n=10$. While it is not common for known multimeric transcription factors to extend beyond tetramers, plotting higher $n$ acknowledges the uncertainty that remains in the field regarding the cooperative mechanics of eukaryotic regulation and particularly of enhancers [60]. Indeed, there are emerging cases of eukaryotic transcription factors functioning as higher-order homomultimeric complexes, such as FOXP3 in regulatory T-cells [91]. It is therefore useful to observe whether such higher-order cooperativities would influence bifurcations in the system.
@@
We find that the bifurcation threshold for the shift between tristable and bistable dynamics is most sensitive to parameter changes at the onset of the expression domain (i.e., at minimum required $n$ and $\bar{a}$). For $n\gtrsim 4$ and $\bar{a}\gtrsim 2$, both parameters are sufficiently large that increasing either has little effect on the threshold probability and corresponding inducer concentration. Given also the existence of homotetrameric transcription factors such as p53 that are crucial to cell fate decisions [54, 46, 57], it is therefore reasonable to choose $n=4$ as representative of dynamics in the highly-cooperative regime, and we do this for the remainder of the paper.
@@
To show more completely how the dynamics evolve with increasing inducer concentration, we now apply the specific parameters $n=4$ and $\bar{a}=2$ to Eqns. 1618. At each inducer concentration, we numerically solve for the fixed point repressor concentrations. We then perform a linear stability analysis on each fixed point. We do this by evaluating the Jacobian which, for our three-gene network modeled by differential equations of the form $d\bar{R}_{i}/dt=f_{i}(\bar{R}_{1},\bar{R}_{2},\bar{R}_{3})$, is a $3\times 3$ matrix with entries
@@
The Jacobian provides a first-order linear approximation of behavior evaluated near a fixed point, with its eigenvectors specifying the primary directions of dynamical motion from the fixed point, and with its eigenvalues reflecting the rates at which perturbations to the fixed point grow or decay along the corresponding eigenvectors. If all eigenvalues have negative real parts, perturbation in any eigenvector direction of expression space will decay back to the fixed point, indicating it to be stable. If any of the eigenvalues have a positive real part, however, perturbation in the corresponding eigenvector direction(s) will grow exponentially and the system will move away from the fixed point, rendering the fixed point unstable.
@@
For the three-gene models we examine, the unstable fixed points are characterized as saddle points. Mathematically, this distinction arises because at least one eigenvalue is negative. Dynamic trajectories local to the fixed point thus move away in some dimensional directions (positive eigenvalues) and approach in others (negative eigenvalues). The number of positive eigenvalues determines the number of dimensional directions of “escape” from the fixed point and thus the “index” of the saddle. If we consider the dynamics of gene expression within a potential landscape, expression levels local to the highest index saddle essentially follow a hierarchical dynamic flow through expression space toward stable states, guided by the presence of saddle points with decreasing index.
@@
Fig. 5 plots the resulting changes in the number of fixed points and their expression levels through bifurcation diagrams, isolating the fixed point trajectories of the $\bar{R}_{1}$ coordinate in panel (A), and the (overlapping) $\bar{R}_{2}$ and $\bar{R}_{3}$ coordinates in panel (B). Three types of fixed points emerge from analysis of the Jacobian. Stable equilibria are denoted in blue, unstable index-1 saddle points in red, and unstable index-2 saddles in yellow.
As the inducer concentration increases, the system shifts across three regimes of decreasing dynamic complexity. Fig. 5(C)-(E) plot fixed points in expression space for a given inducer concentration within each of these regimes. At low inducer concentration, as highlighted in Fig. 5(C), the system is tristable, stabilizing to solely favor expression of one of the three repressors. A sufficiently stochastic system can be perturbed to transition between any pair of these stable points through one of the index-1 saddle points. The system also has an index-2 saddle point at which all repressors are half-maximally expressed. This means that if a system were to have a sufficient concentration of each repressor and initially exist near this index-2 saddle, it would be equally likely for the system to settle at any of the three possible stable states.
@@
At an intermediate inducer concentration, a pair of fixed points annihilate each other and a saddle bifurcation occurs. The resulting intermediate regime, represented in Fig. 5(D), is still tristable, but the three stable states are now only connected by two saddle points. From an initial state with sufficient concentrations of all three repressors, the system can now evolve dynamically in one of two ways: by either falling into a bistable potential well favoring $\bar{R}_{2}$ or $\bar{R}_{3}$ expression, or by following a trajectory directly stabilizing toward $\bar{R}_{1}$-dominant expression. Dynamics are therefore more restricted in this regime than in the low inducer regime, and no longer equally likely to evolve toward any given stable state. Finally, Fig. 5(E) shows fixed point expression levels at a high inducer concentration, where the presence of inducer sufficiently suppresses the activity of $\bar{R}_{1}$ and the system undergoes a saddle bifurcation to collapse down to a bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$-dominant expression.
@@
Fig. 5 thus allows us to determine the inducer concentration thresholds at which dynamic shifts occur, including the shift between tristable and bistable dynamics. Given a distribution of stable state expression levels from data, we could then distinguish between systems in the low inducer regime or in the more dynamically restricted intermediate regime. Fig. 5(B) also highlights that it is only at intermediate inducer concentrations that the system does not necessarily stabilize to a state exclusively favoring a single gene’s expression, and that it is this characteristic that uniquely distinguishes the low vs. intermediate inducer concentration regimes of tristability.
@@
Figure 6: Change in inducer concentration range (shaded regions) of the intermediate dynamic regime for increasing dissociation constant $K_{I}$, overlaid with the corresponding activity probability curves from Fig. 3. Note that regardless of $K_{I}$ parametrization, the intermediate regime is found within the same range of probabilities $0.749\leq p_{\text{act}}(c)\leq 0.867$ (bounded in grey).
@@
Given the relatively narrow concentration range for the intermediate dynamic regime, it is natural to ask how sensitive this window is to changes in allosteric interaction and to changes in the cooperativity, or ultrasensitivity, of expression response to induction. Regarding sensitivity to allostery, altering the ratio of $K_{A}$ and $K_{I}$ for an inducer binding to a repressor shifts the inflection point of the probability curve for activity toward higher inducer concentrations, as previously seen in Fig. 3. We thus observe in Fig. 6 that the log scale range of inducer concentrations within the intermediate dynamic regime does not change but shifts toward increasing value ranges with increasing $K_{I}$. Since altering the ratio of the dissociation constants does not affect the slope of the probability curve at the inflection point, the intermediate dynamic regime (for otherwise fixed values of $m=2$ and $\beta\Delta\epsilon=4$) exists across different $K_{A}/K_{I}$ for the same probability range, $0.749\leq p_{\text{act}}(c)\leq 0.867$.
@@
Figure 7: Change in the regions of multistability for $n\in[3,4]$. Each dynamic region is color-coded as denoted in Fig. 5, with the seven-fixed point tristable regime (purple), five-fixed point tristable regime (light purple), and three-fixed point bistable regime (red). For $n\lesssim 3$, the light purple region illustrates the most complex dynamic regime that is accessible.
@@
The size of the intermediate regime differs, however, with changing cooperativity. Fig.  4(B) indicates that at $\bar{a}=2$, bistability is possible for $n>3$, but the threshold for shifting from tristability to bistability shifts toward lower inducer concentrations for Hill coefficients decreasing from $n<4$. It is thus unclear from Fig. 4 alone whether the range of inducer concentrations allowing the intermediate dynamic regime would expand, contract, or remain the same for a smaller cooperative Hill coefficient. Indeed, Fig. 7 demonstrates how the regimes of multistability evolve in response to varying cooperativity. As the Hill coefficient increases from $n=3$, the inducer concentration at which the system bifurcates from tristable to bistable dynamics rises slightly in the log scale. The threshold separating the two distinct regimes of tristability, however, does significantly change. The inducer concentration range corresponding to the intermediate regime narrows with increasing cooperativity, since this implies a more sensitive response in gene expression to repressor binding, one that emphasizes the bias away from $\bar{R}_{1}$ expression as induction increases. The tristable threshold also begins to approach an apparent asymptote at $n=4$, confirming our earlier finding that increasing cooperativity to $n=4$ has little effect on the dynamics.